Please tell me HOW to work this kind of problem!!

``60(m-15)this is the expression to solve the problems. I sure as heck cant get it to work.

A duplicating machine can produce 60 copies per minute, but the machine takes 15 minutes to warm up before any copies can be made.

1. If this machine produced 2500 copies how long was it on?

I need this problem to be worked step by step. I cant grasp it at all. Please helP!!!





Answer by mananth(16949)

60 copies / minute
Let the number of minutes it worked be m
0.25 hours for warming up.
Total time worked = (m-15) minutes
Total copies produced = 2500
60 copies/minute
2500= 60(m-15)
2500/60 = m-15
m= (2500/60 )+15
m= 56.66 minutes including warmup time